Please help me with this problem
Find the slope
(-12, -8) and (-14, -18)

Answer by Shana-D77(132)

in y = mx + b, "m" os the slope. In physics, slope (aka: rate of change, aka: speed) is found by dividing the change in y (subtract the ys) by the change in x (subtract the xs)
Group 1: (-12, -8) (x1, y1)
Group 2: (-14, -18) (x2, y2)
m = (y2 - y1)/(x2 - x1) (y from group 2 minus y from group 1, etc)
(-18 - -8)/(-14 - -12)
-- = +, so...
(-18 + 8)/(-14 + 12) = -10/-2 = 5!
m = 5